Ask your own question, for FREE!
Mathematics 11 Online
OpenStudy (anonymous):

When a coin is flipped, there are two possible outcomes: heads or tails. If a coin is flipped three times, what is the probability that there will be tails all three times? A.1/8 B.1/3 C.1/2 D.3/2

OpenStudy (raden):

If a coin is flipped three times, all outcomes = {hhh, hht, hth, thh, tth, tht, htt, ttt}, there are 8 space sample. now how many the event of ttt in space sample?

OpenStudy (anonymous):

1/2 probability it will be tails. You want tails all three times so: 1/2*1/2*1/2=1/8

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!